OpenStudy (anonymous):

Find the product: (5x-4y)^2

OpenStudy (saifoo.khan):

(5x-4y)(5x-4y) FOIL

OpenStudy (anonymous):

or (a-b)^2=a^2-2ab+b^2

OpenStudy (anonymous):

so the answer is 25x^2+16y^2

OpenStudy (saifoo.khan):

25 x^2-40 x y+16 y^2

OpenStudy (anonymous):

(5x)(5x)= 25x^2 (5x)(-4y)= -20xy (-4y)(-4y)=16y^2 (-4y)(5x)=-20xy 25x^2-40xy+16y^2

OpenStudy (anonymous):

25x^2-40xy+16y^2